OpenStudy (anonymous):

What are the coordinates of the center of the circle described by the equation (x + 8)2 + y2 = 9?

OpenStudy (campbell_st):

the general form of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] the circle has a centre at (h, k) and radius r so your equation can be written as \[(x + 8)^2 + (y + 0)^2 = 3^2\] hope this helps you find the centre of the circle by matching your equation to the general form.